Saved Drag each label to the appropriate position to Iridicate whether the ltem contalns greater than, equal to, or less than 400 klocalories. The amount of energy required to raise 100,000 ml of water 1 degree Celsius The amount of energy required to raise 400 L of water 1 degree Celsius 40,001 calories 3,999,000 calories The amount of energy required to raise 1 kg of water 1 degree Celsius 400 Calories 400 C 400 с Greater Than 400 Kilocalories Equal To 400 Kilocalories Less Than 400 Kilocalories Reset
Answer #1

1.) 40,0001 calories = Less than 400 Kilocalories

2) The amount of energy required to raise 100,000 ml of water 1 degree celsius : Less than 400 Kilocalories.

3) 3,999,000 = Greater than 400 Kilocalories.

4) The amount of energy required to raise 400 L of water 1 degree celsius = Less than 400 Kilocalories

5).The amount of energy required to raise 1 kg of water 1 degree celsius = Less than 400 Kilocalories

6) 400 Calories = less than 400 Kilocalories

7) 400 C =equal to 400 Kilocalories

8) 400 c = Less than 400 Kilocalories

By dividing calories by 1000 we get Kilocalories, so,

40001 calorie = 40 kilocalorie

3999000 calories = 3999 k calorie

400 calorie = 0.4 kilo calorie

Large calorie or Kilocalorie is written in capital C, while the small calorie is represented by small letter, c

To calculate energy required use the formula Q = mcT

where, m =mass in kilogram, c = specific capacity of water and T = change in temperature of water

so for 1 kg of water at i degree celsius = 1* 4180*1 = 4180 joules

for 100000 ml of water = 100 * 4180*1 = 418000 joules.

for 400 L of water = 0.4 *4180*1 = 1672 joules.

To convert joules to Kilocalorie divide by 4180.